C = 50 + 0.80 (Y – T)
I = 200
G = 100
Find out the equilibrium level of income.
Suppose G increases to 125 what is the new equilibrium level of income?
What level of G is needed to achieve a target income of 200?
1
Expert's answer
2020-05-25T09:25:13-0400

a)"C=50+0.80\u00d7(Y-T)=50+0.80Y-0.8\u00d750=50+0.8Y-40=10+0.8Y;"

"Y=C+I+G;"

"Y=10+0.8Y+200+100;\nY=310+0.8Y;"

"0.2Y=310;"

"Y=1550;"

b)

"G=125;"

"Y=10+0.8Y+200+125; Y=335+0.8Y;"

"0.2Y=335;"

"Y=1675;"

c) "2000=10+0.8*2000+200+G;"

"2000=10+1600+200+G;"

"G=2000-10-1600-200; G=190"

Answer: G equally 190.